### v2

@@ Abstract
In recent years various supervised learning methods that disentangle aleatoric and epistemic uncertainty based on second-order distributions have been proposed. We argue that these methods fail to capture critical components of epistemic uncertainty, particularly due to the mostly neglected component of model bias, but also due to only capturing procedural but not data uncertainty. By using a simulation-based evaluation protocol that encompasses epistemic uncertainty due to both procedural and data uncertainty components, we illustrate that current methods rarely capture the full spectrum of epistemic uncertainty. Through theoretical insights and experimental results on synthetic and real-world data, we find that high model bias can lead to misleadingly low estimates of epistemic uncertainty, and common second-order uncertainty quantification methods for regression seem to blur bias-induced errors into aleatoric estimates, thereby underrepresenting epistemic uncertainty. Our findings underscore that meaningful aleatoric estimates are feasible only if all relevant sources of epistemic uncertainty are properly represented.
keywords:
@@ 1 Introduction
Uncertainty estimation in machine learning has emerged as an important tool to understand the reliability of a model’s predictions. In supervised learning, one often distinguishes between two types of uncertainty: aleatoric and epistemic uncertainty [1, 2]. Assuming that the ground truth data generating process can be described by a conditional probability distribution $p(y|\boldsymbol{x})$, aleatoric uncertainty reflects the inherent randomness of it, and is hence irreducible. Epistemic uncertainty, on the other hand, appears because the unknown $p(y|\boldsymbol{x})$ needs to be estimated using data. As a result, epistemic uncertainty can in principle be reduced by collecting more data, improving model specification, or changing the learning algorithm. While the correct disentanglement [3, 4] of these two types of uncertainty is crucial not only for model interpretability, but also for tasks like out-of-distribution detection [5] and active learning [6], it is far from trivial. Recent works show that using common methods of disentanglement in classification like the information-theoretic decomposition [3] not only is theoretically questionable [7], but can also lead to almost perfectly correlated aleatoric and epistemic uncertainty estimates [8]. In regression, common disentanglement methods like the variance-based decomposition [9, 3, 10] are often used in practice [9, 11, 12], but a thorough theoretical analysis of how well aleatoric and epistemic uncertainty are represented is still missing. Our main goal in this study is therefore to raise three main concerns:
It is questionable which source of epistemic uncertainty existing methods capture. Epistemic uncertainty can arise from multiple sources [13], including distribution shifts between training and test data [14], limited training samples, ill-defined model classes, or procedural randomness in the training algorithm [15]. To represent these uncertainties, most approaches rely on second-order distributions over the parameters of the predictive (first-order) distribution $p(y|\boldsymbol{x})$. This second-order distribution assigns probabilities to all candidate first-order probabilities $\hat{p}(y|\boldsymbol{x})$, thereby quantifying how much the uncertainty estimate can be trusted. Common ways to learn such a second-order representation, include Bayesian neural networks [16, 17], Gaussian processes [18, 19, 20], ensemble models [9] or evidential deep learning [12, 21, 14]. However, it remains unclear which sources of epistemic uncertainty these methods actually capture. Most implicitly focus on the variability of model parameters, while overlooking systematic components such as model bias or hypothesis-space misspecification.
Different sources of epistemic uncertainty make its evaluation troublesome. The ways to evaluate representations of uncertainty are equally vast, where the most common ones are arguably via the performance on downstream tasks [5, 6, 22]. A more theoretical line of work proposes axioms [7] for evaluating second-order distributions, stating several criteria that a second-order distribution has to fulfill in order to faithfully represent epistemic uncertainty. Recently, these axioms were adapted for the regression case [23], highlighting differences between entropy-based and variance-based disentanglement. A more pragmatic approach is the usage of a reference distribution [24] or a gold standard [25], allowing for the relative comparison of second-order distributions. Despite a variety of existing evaluation criteria, we argue that they often only evaluate one specific source of epistemic uncertainty. The aim of out-of-distribution detection methods is typically to capture uncertainty due to a shift between training and test data, while methods under the i.i.d. setting often implicitly assume that any lack of confidence in model parameters arises solely from a limited data sample, neglecting model bias [26] or procedural randomness [15]. Even though the common view that epistemic uncertainty reflects all “lack of knowledge”, practical methods often overlook important contributors, such as biases in the hypothesis space, and the difference between data- and procedural uncertainty [15]. The misinterpretation of which part of the epistemic uncertainty is being captured by common uncertainty quantification methods not only challenges the reliability of their usage for downstream tasks, but also the meaningfulness of the disentanglement between epistemic and aleatoric.
6cm].49
6cm].49
Figure 1: (a): Synthetic experiment where a high model bias is falsely attributed to aleatoric uncertainty: Due to limited data points on the left side of the $x$-axis, and high complexity of the generating function, the underlying epistemic uncertainty there is high. The heteroscedastic Gaussian process model (falsely) predicts it as high aleatoric uncertainty. See Sec. 4 for details. (b): Sources of epistemic uncertainty in machine learning, where $p(y|\boldsymbol{x})$ denotes the true data-generating process. Different sources of epistemic uncertainty can lead to the estimate $\hat{p}(y|\boldsymbol{x})$ being further from the ground truth. See Sec. 2 for further details. Figure adapted from Huang et al. [15].
Our main concern is that correct aleatoric uncertainty estimation is only possible if epistemic uncertainty is sufficiently low. While aleatoric uncertainty arises from the irreducible randomness in the true conditional distribution $p(y|\boldsymbol{x})$, its estimate $\hat{p}(y|\boldsymbol{x})$ depends on how well the model captures the underlying data generating process. Specifically, it is not reliably quantifiable if epistemic uncertainty remains high. If a model has little information about the underlying data generating function, its systematic errors can wrongly be attributed to measurement noise, thereby overestimating aleatoric uncertainty. Conversely, a too low estimate of epistemic uncertainty can lead to false trust in the assessed aleatoric uncertainty. Figure 1(a) shows an exemplary setting where a Gaussian process misclassified high epistemic uncertainty as high aleatoric uncertainty, due to high model bias. Given these concerns, we call for a more fine-grained assessment of disentanglement of aleatoric and epistemic uncertainty. While it is crucial to distinguish between these two sources, one must identify which portion of the epistemic uncertainty a given model truly represents. Furthermore, we argue that one can only make meaningful statements about aleatoric uncertainty once epistemic uncertainty is correctly estimated as sufficiently low.
Contribution. To underline our concerns, we conduct an analysis of epistemic and aleatoric uncertainty estimation for second-order uncertainty representations. Focusing on the regression setting, we examine different sources of epistemic uncertainty, thereby including the often neglected procedural uncertainty, and derive theoretical expressions for them by usage of the classical bias-variance decomposition (Section 2). Subsequently, we discuss how epistemic uncertainty can be evaluated in a correct manner using statistical simulations, and we show with an illustrative example that the underlying model bias can lead to untrustworthy estimates if not taken into account, even if parts of the epistemic uncertainty are correctly estimated (Section 3). Furthermore, by assessing bias, data- and procedural uncertainty conceptually and in synthetic and real data experiments, we illustrate that common second-order uncertainty quantification methods only capture parts of the epistemic uncertainty, often confusing it with aleatoric uncertainty (Section 4).
2 Towards a fine-grained definition of epistemic uncertainty
For the scope of this paper, we define an instance space $\mathcal{X}\subset\mathbb{R}^{D}$, and a label or target space $\mathcal{Y}$, where $\mathcal{Y}\subseteq\mathbb{R}$ for the (univariate) regression, and $\mathcal{Y}=\{1,\dots,K\}$ for the classification case. We further assume having a dataset $\mathcal{D}_{N}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$ of size $N\in\mathbb{N}$, where each $(\boldsymbol{x}_{i},y_{i})$ is an independent and identically distributed (i.i.d.) realization of an unknown probability distribution $P$ on $\mathcal{X}\times\mathcal{Y}$. Let $p(y|\boldsymbol{x})$ be the associated (ground truth) conditional probability to observe label $y$ as an outcome given $\boldsymbol{x}$.
Assuming a parameter space $\Theta$ of the underlying probability distribution, for a given $\boldsymbol{x}\in\mathcal{X}$, we define $\boldsymbol{\theta}(\boldsymbol{x})\in\Theta$ as the parameters of $p(y|\boldsymbol{x})$.
One common representation of uncertainty, which is going to be the focus of this paper, is via a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ over these first-order parameters. Measures such as variance or entropy can then be used to estimate aleatoric and epistemic uncertainty. An estimate $\hat{p}(y|\boldsymbol{x})$ of the predictive distribution can then be obtained by Bayesian model averaging as
@@ 2 Towards a fine-grained definition of epistemic uncertainty
2.1 The difficulty of defining epistemic uncertainty
Unlike aleatoric uncertainty, epistemic uncertainty does not have a universally accepted ground-truth value. Instead, it is inherently context dependent and must always be viewed relative to some benchmark. On the other hand, for different representations one can derive certain logical criteria that the aleatoric and epistemic uncertainty estimates have to fulfill. Empirically, Bengs et al. [27] argued that a second-order distribution should ideally be monotonic with respect to sample size and converge to a Dirac delta distribution if the dataset size goes to infinity. However, we argue that especially these empirical criteria are far from being fully established yet, and miss certain sources of epistemic uncertainty in practical applications, like procedural uncertainty [15], i.e. uncertainty due to randomness in the learning process. To this end, we provide a thorough taxonomy of the different sources of epistemic uncertainty.
As aleatoric uncertainty corresponds to the intrinsic randomness of the data generating process, one can directly associate it with the stochasticity induced by $p(y|\boldsymbol{x})$, using measures such as variance or entropy. Epistemic uncertainty, on the other hand, measures ”everything else” that is in principle reducible by collecting more knowledge, hence all uncertainty with respect to the parameters $\boldsymbol{\theta}$, but also the chosen hypothesis space $\mathcal{H}$. This uncertainty can stem from different sources, and several authors have discussed some of these sources [2, 15, 14]. Unifying those perspectives, we will distinguish in total four different sources of epistemic uncertainty (see Figure 1(b) for an overview). Having a loss function $\ell:\mathbb{P}(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ with associated expected loss $\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(\cdot,y)\big]$, a first important subdivision into three components can be made as follows:
  • Model uncertainty: arises when the hypothesis space $\mathcal{H}$ does not contain the true underlying distribution, i.e. $p(y|\boldsymbol{x})\mathrel{\vtop{\halign{#\cr\hfil$\in$\hfil\cr\hfil$/$\hfil\cr}}}\mathcal{H}$. In Figure 1(b) the notation $\hat{p}(y|\boldsymbol{x})$ is used to denote the Bayes-optimal model for a given hypothesis space. It can be formally defined as follows:
    $\hat{p}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{y|\boldsymbol{x}}\ell(\hat{p}(y|\boldsymbol{x}),y).$
  • Estimation uncertainty: quantifies how well the learning algorithm can approximate $\hat{p}(y|\boldsymbol{x})$, which is the best possible model given the hypothesis space.
  • Distributional uncertainty: is due to a shift in distribution between training and test data [14], respectively, and their associated conditional distributions $p(y|\boldsymbol{x})$ and $\tilde{p}(y|\boldsymbol{x})$.
Following recent work of Huang et al. [15], we further divide estimation uncertainty into data uncertainty and procedural uncertainty. To this end, we define $\mathcal{D}_{N}\sim P^{N}$ as the random variable reflecting the data uncertainty, where $P^{N}$ is the distribution over all datasets of size $N$. Similarly, we define $\gamma\sim P^{\gamma}$ as the random variable reflecting all procedural parameters (e.g. random weight initialization in Deep Ensembles), where $P^{\gamma}$ denotes the distribution over which these parameters are sampled. For a fixed dataset $\mathcal{D}_{N}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$, data uncertainty can then be formally defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and
$\hat{p}_{N}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{\gamma}\big[\frac{1}{N}\sum_{i=1}^{N}\ell(\hat{p}_{N,\gamma}(y|\boldsymbol{x}_{i}),y_{i})\big]\,$,
where $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ denotes a model trained on a given dataset of size $N$ and procedural parameters $\gamma$. Similarly, for a fixed set of procedural parameters, procedural uncertainty can be formally defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and
$\hat{p}_{\gamma}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{\mathcal{D}_{N}}\big[\frac{1}{N}\sum_{i=1}^{N}\ell(\hat{p}_{N,\gamma}(y|\boldsymbol{x}_{i}),y_{i})\big].$
Using this notation, estimation uncertainty is then defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$, as visualized in Figure 1(b).
2.2 Uncertainty from a generalized bias-variance perspective
From a statistician’s viewpoint, the classical bias-variance decomposition provides further insight into the interplay between estimation uncertainty and aleatoric uncertainty [20, 13]. While it is originally defined for the mean squared error as a loss function, it can also be generalized to any Bregman divergence [28, 29]. With $\ell:\mathbb{P}(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ being a proper scoring rule, for a given estimate $\hat{p}(y|\boldsymbol{x})$ of the conditional distribution, one can decompose the point-wise risk as
$\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(\hat{p}(y|\boldsymbol{x}),y)\big]$ $=\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(p(y|\boldsymbol{x}),y)\big]+\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(p(y|\boldsymbol{x}),y)-\ell(\hat{p}(y|\boldsymbol{x}),y)\big]$
@@ 2.2 Uncertainty from a generalized bias-variance perspective
$=\underbrace{\mathbb{H}_{\ell}\big[p(y|\boldsymbol{x})\big]}_{\text{aleatoric uncertainty}}+\underbrace{D_{\ell}\big(p(y|\boldsymbol{x}),\hat{p}(y|\boldsymbol{x})\big)}_{\text{epistemic uncertainty}}$ (2)
where $\mathbb{H}_{\ell}$ is the generalized Bayes risk of $\ell$ and $D_{\ell}$ is its associated divergence. For the empirical estimate, let $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ be the predictive distribution obtained with data uncertainty $D_{N}\sim P^{N}$ and procedural uncertainty $\gamma\sim P^{\gamma}$. Then the pointwise expected loss becomes [30]
@@
$\mathbb{E}_{y|\boldsymbol{x}}\big[\mathbb{E}_{\mathcal{D}_{N},\gamma}[\ell(\hat{p}_{N,\gamma}(y|\boldsymbol{x}),y)]\big]=$ $\mathbb{H}_{\ell}\big[p(y|\boldsymbol{x})\big]+\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(p(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x}))\big]$
$=$ $\underbrace{\mathbb{H}_{\ell}\big[p(y|\boldsymbol{x})\big]}_{\text{aleatoric unc.}}+\underbrace{D_{\ell}\big(p(y|\boldsymbol{x}),\bar{p}^{*}(y|\boldsymbol{x})\big)}_{\text{generalized bias}}$

### v3

@@ Abstract
Identifying and disentangling sources of predictive uncertainty is essential for trustworthy supervised learning. We argue that widely used second-order methods that disentangle aleatoric and epistemic uncertainty are fundamentally incomplete. First, we show that unaccounted bias contaminates uncertainty estimates by overestimating aleatoric (data-related) uncertainty and underestimating the epistemic (model-related) counterpart, leading to incorrect uncertainty quantification. Second, we demonstrate that existing methods capture only partial contributions to the variance-driven part of epistemic uncertainty; different approaches account for different variance sources, yielding estimates that are incomplete and difficult to interpret. Together, these results highlight that current epistemic uncertainty estimates can only be used in safety-critical and high-stakes decision-making when limitations are fully understood by end users and acknowledged by AI developers.
@@ 1 Introduction
Figure 1: Conceptual illustration of our two main arguments for the regression case. Left: When high model bias is present, we show that aleatoric uncertainty overestimates the ground truth noise. (Sec. 3.1). Right: Many second-order methods capture only procedural and not data uncertainty (Sec. 3.2), thereby underestimating the variance part of the epistemic uncertainty.
@@
Uncertainty estimation in machine learning is essential for assessing the reliability of model predictions. In recent supervised learning research, one often distinguishes between two types of uncertainty: aleatoric and epistemic uncertainty (Kendall and Gal, 2017; Hüllermeier and Waegeman, 2021). Assuming that the ground truth data-generating process can be described by a conditional probability distribution $p(y|\boldsymbol{x})$, aleatoric uncertainty reflects the inherent randomness of this distribution, and is hence irreducible. Epistemic uncertainty, on the other hand, appears because the unknown $p(y|\boldsymbol{x})$ needs to be estimated using data. As a result, epistemic uncertainty can be reduced by collecting more data, improving model specification, or changing the learning algorithm. While the correct disentanglement (Depeweg et al., ; Valdenegro-Toro and Mori, 2022) of these two types of uncertainty is crucial for model interpretability, and for downstream tasks such as out-of-distribution detection (Hendrycks and Gimpel, 2017) and active learning (Gal et al., 2017), it is far from trivial.
One principled way of representing epistemic uncertainty is through a second-order distribution, that is, a distribution over possible predictive distributions (Bengs et al., 2022). In this work, we focus on second-order distributions because many existing uncertainty quantification approaches can be interpreted within this framework, including Bayesian methods, ensemble methods, and deterministic approaches (Gawlikowski et al., 2023; He et al., 2025). In Bayesian models, uncertainty is represented by a posterior distribution over model weights, which induces a distribution over a classical first-order distribution (e.g. a distribution over mean and variance of a Gaussian in a regression setting). In posterior predictive sampling, one takes multiple samples of the weights from the posterior distribution and computes the corresponding second-order distribution (Gal and Ghahramani, 2016; Denker and LeCun, 1990; Neal, 2012; Blundell et al., 2015). Ensemble methods similarly induce a distribution over first-order distributions by aggregating outputs from multiple independently trained models (Lakshminarayanan et al., 2017; Ovadia et al., 2019). Finally, some approaches aim to model epistemic uncertainty more directly by learning the parameters of the second-order distribution using specific loss functions (Charpentier et al., 2021; Amini et al., 2020; Sensoy et al., 2018; Malinin and Gales, 2018b). Viewed through this lens, these approaches provide a common framework for disentangling aleatoric and epistemic uncertainty.
Under a second-order representation, uncertainty can be disentangled in different ways depending on the learning task. In classification, uncertainty is commonly analyzed using information-theoretic decompositions, which separate total, aleatoric, and epistemic uncertainty (Depeweg et al., ). For regression, analogous variance-based decompositions are typically employed (Lakshminarayanan et al., 2017; Depeweg et al., ; Sale et al., 2023). However, a growing body of critical work argues that such second-order decompositions are not theoretically well-founded and can be difficult to interpret in practice. For evidential second-order risk minimization, recent analyses highlight identifiability and convergence pathologies, showing that the resulting epistemic uncertainty estimates are tightly controlled by regularization, making them only meaningful in a relative sense (Bengs et al., 2022; Meinert et al., 2023; Jürgens et al., 2024). Complementarily, in Bayesian deep learning, (Fellaji and Pennerath, 2024) report failure modes in which epistemic uncertainty can collapse as model size increases, contrary to what is expected. More broadly, (Kirchhof et al., 2025) emphasizes that aleatoric/epistemic definitions and decomposition principles can be mutually inconsistent, and that practical estimators may become strongly correlated, challenging the premise that these components are meaningfully disentangled. In the same sense, recent works show that these decompositions are theoretically dubious (Wimmer et al., 2023; Smith et al., 2024), and may produce highly correlated aleatoric and epistemic uncertainty estimates (Mucsányi et al., 2024), questioning their ability to disentangle different sources of uncertainty meaningfully.
Our position. The most prominent recent papers that propose uncertainty disentanglement methods have collectively received almost 40,000 citations [^1] Number obtained from Google Scholar. Check Appendix A for details., and their usage in safety-critical applications is increasing. This relevance, together with the recent critical work motivate a more fundamental reconsideration of what epistemic uncertainty is intended to represent and quantify in predictive models. Accordingly, we argue that recent and highly-cited epistemic uncertainty estimation methods are fundamentally incomplete. To substantiate this claim, we proceed as follows. In Section 2, we review and systematize the different sources of epistemic uncertainty identified in the literature, distinguishing between estimation uncertainty (variance-related), model uncertainty (bias-related), and distributional uncertainty. In Section 3.1 we adopt the bias–variance decomposition of Bregman divergences (Hastie et al., 2009; Gruber et al., 2023) and demonstrate that the bias component is as central to epistemic uncertainty as the variance component. Within this framework, we show that bias induced by the regularization mechanisms commonly used in neural networks can substantially distort uncertainty estimates, even when the true data-generating function lies within the model’s hypothesis space. In Section 3.2, we turn to the variance component and demonstrate that existing uncertainty quantification methods fail to capture it in its entirety; instead, different approaches account for different variance contributions, resulting in estimates that are inherently incomplete and difficult to interpret. Finally, in Section 4 we discuss alternative perspectives on epistemic uncertainty proposed in the literature and relate them to our work.
2 Setup and preliminaries
Supervised learning setup.
We consider an instance space $\mathcal{X}\subset\mathbb{R}^{D}$ and label space $\mathcal{Y}$, where $\mathcal{Y}\subseteq\mathbb{R}$ for regression and $\mathcal{Y}=\{1,\ldots,K\}$ for classification. We assume a dataset $\mathcal{D}_{N}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$ of i.i.d. samples from an unknown joint distribution $P$ on $\mathcal{X}\times\mathcal{Y}$, with ground-truth conditional distribution $p(y|\boldsymbol{x})$. As a loss function $\ell:\mathcal{P}(\mathcal{Y})\times\mathcal{Y}\to\mathbb{R}$ we assume a strictly proper scoring rule (Gneiting and Raftery, 2007) which uniquely minimizes the expected loss when the prediction equals the ground truth. Every strictly proper scoring rule induces a convex entropy $\mathbb{H}_{\ell}$ and a Bregman divergence $D_{\ell}$ such that the pointwise expected score gap equals the divergence:
$D_{\ell}(p(\cdot|\boldsymbol{x}),\hat{p}(\cdot|\boldsymbol{x})):=\mathbb{E}_{y|\boldsymbol{x}}[\ell(\hat{p}(\cdot|\boldsymbol{x}),y)]-\mathbb{E}_{y|\boldsymbol{x}}[\ell(p(\cdot|\boldsymbol{x}),y)],$
with entropy $\mathbb{H}_{\ell}(p(\cdot|\boldsymbol{x})):=\mathbb{E}_{y|\boldsymbol{x}}[\ell(p(\cdot|\boldsymbol{x}),y)]$ (Banerjee et al., 2005; Gneiting and Raftery, 2007). For a hypothesis space $\mathcal{H}:=\{p_{\boldsymbol{\theta}}:\mathcal{X}\to\mathcal{P}(\mathcal{Y})|\boldsymbol{\theta}\in\Theta\}$, of possible predictive distributions parametrized by $\boldsymbol{\theta}\in\Theta$, the risk of an estimator $\hat{p}\in\mathcal{H}$ integrates the pointwise expected loss over $\boldsymbol{x}$:
$\mathcal{L}(\hat{p}):=\mathbb{E}_{\boldsymbol{x}}\left[\mathbb{E}_{y|\boldsymbol{x}}[\ell(\hat{p}(\cdot|\boldsymbol{x}),y)]\right]=\mathbb{E}_{(\boldsymbol{x},y)}[\ell(\hat{p}(\cdot|\boldsymbol{x}),y)].$
By strict properness, the unconstrained risk minimizer is $p(y|\boldsymbol{x})$ itself. The Bayes-optimal estimator within $\mathcal{H}$ is
[…]
Second-order distributions. Following common literature (Bengs et al., 2022; Hüllermeier and Waegeman, 2021; Gawlikowski et al., 2023; Li et al., 2025), we represent uncertainty via a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ over parameters $\theta\in\Theta$ of the (first-order) distribution. The predictive distribution can then obtained by model averaging:
@@ 2 Setup and preliminaries
Existing measures that decompose the uncertainty of $q(\boldsymbol{\theta}|\boldsymbol{x})$ into an aleatoric and epistemic component focus on entropy (Depeweg et al., ), distance to reference sets (Sale et al., 2023), or proper scoring rules (Hofman et al., 2024; Fishkov et al., 2025). In regression, the variance-based decomposition (Depeweg et al., ; Sale et al., 2023) is most commonly used. In this approach, the variance of the prediction $\hat{y}$, obtained via model averaging as in Eqn. 2, is decomposed into an aleatoric and an epistemic part:
@@
$\text{Var}(\hat{y}|\boldsymbol{x})=\underbrace{\mathbb{E}_{\boldsymbol{\theta}}\big[\text{Var}(\hat{y}|\boldsymbol{x},\boldsymbol{\theta})\big]}_{\text{ aleatoric estimate}}+\underbrace{\text{Var}_{\boldsymbol{\theta}}\big(\mathbb{E}[\hat{y}|\boldsymbol{x},\boldsymbol{\theta}]\big)}_{\text{epistemic estimate}}.$ (3)
Sources of epistemic uncertainty.
Figure 2: Sources of epistemic uncertainty in machine learning, where $p(y|\boldsymbol{x})$ denotes the true data-generating process. Different sources of epistemic uncertainty can lead to the estimate $\hat{p}(y|\boldsymbol{x})$ being further from the ground truth. See Sec. 2 for further details. Figure inspired from Huang et al. (2023)
As aleatoric uncertainty corresponds to the intrinsic randomness of the data-generating process, one can directly associate it with the stochasticity induced by the ground truth conditional distribution, using measures such as variance or entropy. Epistemic uncertainty, on the other hand, measures the uncertainty about the ground truth, i.e., the expected deviation from it. It is reducible by collecting more knowledge, hence decreasing the uncertainty about the parameters $\boldsymbol{\theta}$, but also the chosen hypothesis space $\mathcal{H}$. This uncertainty can stem from different sources, which several authors discuss (Hüllermeier and Waegeman, 2021; Huang et al., 2023; Malinin and Gales, 2018b). Unifying those perspectives, we will distinguish in total five distinct sources of epistemic uncertainty (see Figure 2): approximation error, estimation bias, data uncertainty, procedural uncertainty and distributional uncertainty.
Following (Huang et al., 2023), we define $\mathcal{D}_{N}\sim P^{N}$ as the random variable reflecting the data uncertainty, where $P^{N}$ is the distribution over all datasets of size $N$. Similarly, let $\gamma\sim P^{\gamma}$ be the random variable reflecting all procedural uncertainty, which arises from the training process from a single model (due to e.g. random weight initialization or data ordering) and is present even with infinite data. $P^{\gamma}$ denotes the distribution over which the procedural parameters are sampled. The estimator $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ depends on both types of uncertainties. We define the dual mean as the Bregman centroid (Banerjee et al., 2005; Pfau, 2013):
@@
$\bar{p}^{*}(y|\boldsymbol{x}):=\operatorname*{\arg\min}_{\tilde{p}\in\mathcal{P}(\mathcal{Y})}\mathbb{E}_{\mathcal{D}_{N},\gamma}\left[D_{\ell}(\tilde{p},\hat{p}_{N,\gamma}(\cdot|\boldsymbol{x}))\right].$ (4)
Note that $\bar{p}^{*}$ does not have to lie in $\mathcal{H}$, as depicted in Figure 2 For example, when $\mathcal{H}$ only consists of decision stumps, the dual mean might be a more complex decision tree. One can now divide the epistemic uncertainty into:
Model uncertainty: discrepancy between $\bar{p}^{*}(y|\boldsymbol{x})$ and $p(y|\boldsymbol{x})$. It can be further decomposed into estimation bias, the systematic deviation of $\bar{p}^{*}(y|\boldsymbol{x})$ from the Bayes estimator $\hat{p}^{*}(y|\boldsymbol{x})$ (Eq. 1) and approximation error, which is the difference between $\hat{p}^{*}(y|\boldsymbol{x})$ and $p(y|\boldsymbol{x})$ (Brown and Ali, 2024):
$\mathbb{E}_{(\boldsymbol{x},y)}[D_{\ell}(p(\cdot|\boldsymbol{x}),\bar{p}^{*}(\cdot|\boldsymbol{x}))]$ (5)
$=\underbrace{\mathcal{L}(\hat{p}^{*}(y|\boldsymbol{x}))-\mathcal{L}(p(y|\boldsymbol{x}))}_{\text{approximation error}}+\underbrace{\mathcal{L}(\bar{p}^{*}(y|\boldsymbol{x}))-\mathcal{L}(\hat{p}^{*}(y|\boldsymbol{x}))}_{\text{estimation bias}}.$
Estimation uncertainty: variation of $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ around the dual mean $\bar{p}^{*}(y|\boldsymbol{x})$. It further decomposes into data and procedural uncertainty. For a fixed dataset $\mathcal{D}_{N}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$, data uncertainty can be defined as the difference between $\bar{p}^{*}(y|\boldsymbol{x})$ and
$\hat{p}_{N}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}\in\mathcal{H}}\mathbb{E}_{\gamma}\big[\frac{1}{N}\sum_{i=1}^{N}\ell(\hat{p}_{N,\gamma}(\cdot|\boldsymbol{x}_{i}),y_{i})\big],$
where the expected empirical loss is taken over the procedural parameters $\gamma$. Similarly, for fixed procedural parameters $\gamma$, procedural uncertainty can be defined as the difference between $\bar{p}^{*}(y|\boldsymbol{x})$ and
$\hat{p}_{\gamma}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}\in\mathcal{H}}\mathbb{E}_{\mathcal{D}_{N}}\big[\frac{1}{N}\sum_{i=1}^{N}\ell(\hat{p}_{N,\gamma}(\cdot|\boldsymbol{x}_{i}),y_{i})\big].$
Using this notation, estimation uncertainty is the difference between $\bar{p}^{*}(y|\boldsymbol{x})$ and $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$, as visualized in Figure 2.
Distributional uncertainty: distribution shift between training and test data (Malinin and Gales, 2018b), respectively, and their associated conditional distributions $p(y|\boldsymbol{x})$ and $\tilde{p}(y|\boldsymbol{x})$.
3 Epistemic uncertainty estimation methods are fundamentally incomplete
In this section we argue that current approaches to estimating aleatoric and epistemic uncertainty are fundamentally incomplete. To our opinion, this incompleteness has two main reasons:
  • First, estimation bias and approximation error both distort epistemic and aleatoric uncertainty estimates. Since second-order representations do not have access to the ground truth conditional class distribution, it is not possible to incorporate model uncertainty into epistemic uncertainty estimates. Consequently, a complete representation of epistemic uncertainty is fundamentally unattainable (Section 3.1).
  • Second, existing estimators are also incomplete because they do not represent data uncertainty and procedural uncertainty at the same time. From a bias-variance perspective, these two types of uncertainty are closely related to the variance of predictions (see Section 3.2).
To illustrate our arguments, we will focus on the estimation of uncertainty in a regression setting. Much of the existing critical literature on uncertainty disentanglement has concentrated on the classification setting, but the regression setting allows us to present the core ideas more transparently, while remaining complementary to prior work. We consider a data-generating process such that $y_{i}=f(\boldsymbol{x}_{i})+\epsilon_{i}$ and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(\boldsymbol{x}_{i}))$. Here, $\epsilon_{i}$ is data-dependent (heteroscedastic) noise, and $f$ is a deterministic function, such that $p(\cdot|\boldsymbol{x})$ forms the density of a Gaussian distribution with mean $f(\boldsymbol{x})$ and variance $\sigma^{2}(\boldsymbol{x})$.
For illustrations we make use of the synthetic regression setting of Rossellini et al. (2024). A dataset $\mathcal{D}_{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$ is generated where $x_{i}\sim Beta(1.2,0.5)$ and
$y_{i}=\sin\left(\frac{1}{5(x_{i}+0.16)^{3}}\right)+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(x_{i})),$ (6)
with $\sigma^{2}(x)=x^{4}$. This data-generating process is chosen because it represents a problem with two main regions. The left side of the data domain ($x<0.4$) has low true aleatoric uncertainty and a difficult-to-learn function, which is expected to induce a high epistemic uncertainty. The right side of the data domain ($x\geq 0.4$) has high true aleatoric uncertainty and a simple function (which is expected to induce a low epistemic uncertainty). This one-dimensional setting is ideal for visually illustrating our arguments regarding uncertainty disentanglement (Figures 3 and 4). As an example, we will illustrate the behavior of Deep Ensembles, considering a simple multi-layer perceptron that outputs $\hat{f}(x)$ and $\hat{\sigma}^{2}(x)$ as the mean and variance of the Gaussian distribution at $x$. Predictions coming from different parametrizations of the neural network are then aggregated to obtain a second-order distribution. In Appendix C we provide similar illustrations for a real-world regression dataset and for other second-order models.
3.1 Bias distorts uncertainty estimations.
From a statistician’s viewpoint, the classical bias-variance decomposition provides further insight into the interplay between estimation uncertainty and aleatoric uncertainty (Hastie et al., 2009; Gruber et al., 2023). While it is originally defined for the mean squared error as a loss function, it can also be generalized to any proper scoring rule (Gruber and Buettner, 2023; Pfau, 2013). As a small further extension, we incorporate data and procedural components into the decomposition:
[…]